OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the given circle. I know how to do this, but I'm confused what the x or y is equal to. 1, 0, x, or y? (x)^2 + (y)^2 = 16 center at point ( , ) radius = 4

OpenStudy (btaylor):

the general form of a circle with center at (h,k) and radius r is:\[(x-h)^2 + (y-k)^2 = r^2\]In your problem, h = 0 and k = 0. So the center is at the origin, (0,0)